A boy throws a baseball onto a roof and it rolls back down and off the roof with a speed of 3.80 m/s. If the roof is pitched at 24.0° below the horizon and the roof edge is 3.30 m above the ground, find the time the baseball spends in the air and the horizontal distance from the roof edge to the point where the baseball lands on the ground.


1
Expert's answer
2022-05-22T14:37:16-0400
"y=y_0-v_0\sin\theta t-gt^2/2\\
y=3.30-1.55 t-4.9t^2"

At the grounf level

"y=0\\
3.30-1.55 t-4.9t^2=0"

The time of flight

"t=0.678\:\rm s"

The horizontal distance

"R=v_0\cos\theta t\\
=3.80\cos 24.0^\circ*0.678=2.35\:\rm m"